OpenStudy (anonymous):

Evaluate: 3√8x^3 - 1 = 2x-1

OpenStudy (whpalmer4):

Is that \[\sqrt[3]{8x^3} -1 = 2x -1\] or \[\sqrt[3]{8x^3-1} = 2x - 1\]

OpenStudy (whpalmer4):

Ah, must be the latter. Try cubing both sides...

OpenStudy (anonymous):

it's actually \[4\sqrt{3-8x ^{2}} = 2x-1\]

OpenStudy (anonymous):

sorry, i had the wrong equation!

OpenStudy (whpalmer4):

Same approach :-)

OpenStudy (anonymous):

i multiplied both sides by 4 and ended up with 3-8x^2 = 16x^4

OpenStudy (anonymous):

how should i factor this to find x?

OpenStudy (whpalmer4):

How did multiplying both sides by 4 get you that? And is that 4 * the square root, or the 4th root?

OpenStudy (anonymous):

it's the 4th root

OpenStudy (anonymous):

multiplying by 4 cancels out the 4th root

OpenStudy (whpalmer4):

Okay, to write nth root, you type \ ( \sqrt[n]{stuff} \ )

OpenStudy (anonymous):

\[\sqrt[4]{3-8x ^{2}} = 2x\]

OpenStudy (anonymous):

then i got the 3-8x\[3-8x ^{2} = 16x^{4}\]

OpenStudy (anonymous):

or should i have ended up with 2x instead of 16x^{4}

OpenStudy (whpalmer4):

multiplying by 4 cancels the 4th root? huh?

OpenStudy (anonymous):

yeah... o.o

OpenStudy (anonymous):

okay nevermind... i'll still give you the "best response" medal though. thanks for your time

OpenStudy (whpalmer4):

Sorry, having some browser problems. You didn't multiply by 4, you raised each side to the 4th power, big difference! \[(\sqrt[4]{3-8x^2})^4 = (2x)^4\]\[3-8x^2=2^4x^4 = 16x^4\]\[16x^4+8x^2+3 = 0\] Now how do we factor that?

OpenStudy (anonymous):

yes! that's exactly where i'm stuck. i have no idea how to factor this kind of trinomial

OpenStudy (anonymous):

by the way shouldn't it be \[16x ^{4} + 8x ^{2} - 3 = 0\], not +3?

OpenStudy (whpalmer4):

Well, this looks like a quadratic, except instead of \(ax^2+bx+c = 0\) we've got \(ax^4 + bx^2 + c = 0\) Let's try a substitution to make it look like a quadratic. We'll do \(u = x^2\) \[16u^2 +8u -3 = 0\] (yes, good catch on the sign error!) That we can solve by any of the usual methods for quadratics. Then we plug \(x^2\) in wherever we see \(u\) in the solution.

OpenStudy (whpalmer4):

Yes, much nicer roots with the -3 vs. the +3 :-)

OpenStudy (anonymous):

but how do i solve this once i've plug in x ^{2} ?

OpenStudy (anonymous):

\[16x(x ^{2})^{2} + 8(x ^{2}) - 3 = 0\] ?? what do i do next?

OpenStudy (anonymous):

i can't find any factors that add up to 8 and have product of -3

OpenStudy (whpalmer4):

No, now you solve \[16u^2 + 8u - 3 = 0\]What are the solutions?

OpenStudy (whpalmer4):

Use the quadratic formula

OpenStudy (anonymous):

so this equation with the u's does not factor?

OpenStudy (whpalmer4):

Oh, you can factor it. To factor that, you multiply 16 and -3 = -48. Now find a pair of factors of -48 that add to 8. -4 and 12 would work. \[(4u -1 )(4u+3 ) = 16u^2+12u-4u-3 = 0\] Pair of real roots:\[4x^2 -1 = 0\] Pair of complex roots:\[4x^2+3=0\]

OpenStudy (whpalmer4):

I just didn't want to explain how to factor it :-)

OpenStudy (anonymous):

haha! okay, well thank you so much for your help :)

OpenStudy (whpalmer4):

Can you find the roots from the remaining two equations?

OpenStudy (anonymous):

i think so.

OpenStudy (whpalmer4):

I'll check your work if you want when you've got them.

OpenStudy (anonymous):

i got 4x^2 +1, not 4x^2 - 1. also, how do i find x? i know that the correct answer is that x=1/2 (from the back of my math book) but i'm not sure why.

OpenStudy (anonymous):

based on \[4x ^{2} +1 = 0\] ....how do i find x? i can't find the square root of \[-\frac{ 1 }{ 4 }\]

OpenStudy (anonymous):

OH nevermind. you were right, it's negative 1. okay, i got 1/2 now. THANK YOU!!!

OpenStudy (whpalmer4):

Well, it's both 1/2 and -1/2, because (-1/2)^2 = 1/4...

OpenStudy (whpalmer4):

Now how about the (4x^2 +3) = 0? You've got a pair of roots there, too.

OpenStudy (whpalmer4):

Hint: remember that i^2 = -1, so that can be written as (4x^2 - (-1)(3)) = 0 or 4x^2 - 3i^2 = 0 or 4x^2 = 3i^2

OpenStudy (whpalmer4):

@shawnomnom whenever you square or cube or whatever both sides of the equation, you need to check your answers to make sure that they actually solve the original equation. often, you will get some solutions that do not work.